# Spherical Knot Mosaics

### v1

@@ 5 Bounds on Spherical Knot Mosaic Invariants
    Each tile in the sequence is one of $T_{5},T_{6},T_{9}$ or $T_{10}$.
@@
    Consecutive tiles in the sequence are edge-adjacent and have connection points on opposite edges so that the sequence forms a continuous, non-branching closed path with no turns or gaps, but possibly under or over crossings.
@@
    Removing the tiles which form the mosaic belt disconnects the spherical mosaic.
@@
If $SM_{K}$ is a spherical knot $n$-mosaic for a knot, $K$, then maximum number of $T_{9}$ or $T_{10}$ tiles of $SM_{K}$ is $6n^{2}-3n+1$.
@@
We will start with a spherical $n$-mosaic containing only crossing tiles, $T_{9}$ or $T_{10}$. There are $6n^{2}$ total tiles. This is not a spherical knot $n$-mosaic, but is a spherical link $n$-mosaic with exactly $3n$ connected components each forming a spherical mosaic belt. We will begin eliminating crossings one at a time until there is only one connected component.
Choose any spherical mosaic belt. Now choose any tile on that belt and replace it to a non-crossing tile with four connection points, $T_{7}$ or $T_{8}$. This change reduces the number of connected components in our spherical link $n$-mosaic by one and means the tile opposite the one we changed now represents a crossing of a single component instead of two disjoint components.
We can then iterate this process by choosing another crossing tile that is part of our larger connected component which does not represent the component crossing itself and changing it to a non-crossing tile with four connection points, this again reduces the number of connected components by one. We can do this until there is only one connected component and thus we have a spherical knot $n$-mosaic for some alternating knot $K$ with $6n^{2}-(3n-1)$ crossing tiles as there were originally $6n^{2}$ crossing tiles and we removed 3n13n-1 of them. This is the most crossings possible for a spherical knot $n$-mosaic. ∎
@@
We modify the proof of Theorem 5.6 such that the original tiling with $6n^{2}$ crossing tiles alternates. That is, for each component every under crossing is preceded by and followed by an over crossing and vice versa. Now notice that replacing any crossing tile involving two components with $T_{7}$ or $T_{8}$ preserves the alternating nature of the link diagram, thus after replacing 3n13n-1 crossing tiles we are left with an alternating knot diagram with $6n^{2}-3n+1$ crossings. Note this may not be a reduced knot diagram so we do not necessarily find that this is the crossing number, but instead an upper bound for the crossing number of our knot. ∎
Figure 6: An example of a 22- spherical knot mosaic which has the maximal number of crossing tiles of $6\cdot 2^{2}-3\cdot 2+1=19$.
@@
From Corollary 5.6.2 we have that
Proposition 5.7.
@@
Similarly from Corollary 5.6.2 we have that
Proposition 5.8.
Let $SM_{K}$ be a spherical knot $2$-mosaic for some knot $K$, then $19\geq c(K)$.
which further implies
Corollary 5.8.1.
Let $K$ be a knot such that $c(K)\geq 20$, then $sm(K)\geq 3$.
There are several more questions spherical knot mosaics gives rise to which the authors plan to explore in subsequent papers.
@@ Acknowledgments
The authors would like to thank Jennifer Schultens for her encouragement and advice throughout this project.
@@ Appendix A Planar Representation Examples
Planar representations of the first 35 unoriented prime knots as spherical knot mosaics. Due to Corollary 5.6.2, each of these planar representations realize their knot type’s spherical mosaic number.
[Uncaptioned image][Uncaptioned image][Uncaptioned image]

### v2

@@ 5 Bounds on Spherical Knot Mosaic Invariants
    Each tile in the sequence has exactly two connecting edges to adjacent tiles.
@@
    Consecutive tiles in the sequence share an edge and connected through opposite edges so that the sequence forms a continuous, non-branching path with no turns or gaps.
@@
    The complement of the mosaic belt is disconnected.
Figure 12: The six distinct mosaic belts of the empty spherical 2-mosaic
@@
If $SM_{K}$ is a spherical knot $n$-mosaic for a knot, $K$, then the maximum number of crossing tiles in $SM_{K}$ is $6n^{2}-3n+1$.
@@
We will start with a spherical $n$-mosaic containing only crossing tiles, so it is entirely tiled with $T_{9}$ or $T_{10}$. There are $6n^{2}$ total tiles. This is not a spherical knot $n$-mosaic, but instead is a spherical link $n$-mosaic with exactly $3n$ connected components as each connected component corresponds to one of the $3n$ mosaic belts. We will replace each crossing tile one at a time with $T_{7}$ or $T_{8}$ until there is only one connected component remaining.
Choose any mosaic belt. Now choose any tile on that belt and replace it with a non-crossing tile with four connection points, either $T_{7}$ or $T_{8}$. This change reduces the number of connected components in our spherical link $n$-mosaic by one and means the tile opposite the one we changed now represents a crossing of a single component instead of two disjoint components.
We can then iterate this process by choosing another crossing tile that is part of our larger connected component which does not represent the component crossing itself and changing it to a non-crossing tile with four connection points, this again reduces the number of connected components by one. We can do this until there is only one connected component and thus we have a spherical knot $n$-mosaic for some knot with $6n^{2}-(3n-1)$ crossing tiles as there were originally $6n^{2}$ crossing tiles and we removed 3n13n-1 of them. This is therefore the most crossings possible for a spherical knot $n$-mosaic. ∎
Remark 5.7.
Theorem 5.6 is an equivalent result to [8, Theorem 3.2 p. 634]
@@
We modify the proof of Theorem 5.6 such that the original tiling with $6n^{2}$ crossing tiles alternates. That is, for each component every under crossing is preceded by and followed by an over crossing and vice versa. Now notice that replacing any crossing tile involving two components with $T_{7}$ or $T_{8}$ preserves the alternating nature of the diagram, thus after replacing 3n13n-1 crossing tiles we are left with an alternating knot diagram with $6n^{2}-3n+1$ crossings. Note this may not be a reduced knot diagram so we do not necessarily find that this is the crossing number, but instead an upper bound for the crossing number of our knot. ∎
Remark 5.8.
One can find a stronger result in [8, Theorem 3.3 p. 636]. They show there will always exist a knot such that $sm(K)=n$ and $c(K)=6n^{2}-3n+1$.
Figure 13: An example of a spherical knot $2$-mosaic which has the maximum possible number of crossing tiles, $6\cdot 2^{2}-3\cdot 2+1=19$
@@
From Corollary 5.8.1 we have the following two results.
Corollary 5.8.2.
@@
We summarize these and further results of Corollary 5.8.1 in Table 3.
If $sm(K)$, then $c(K)$
1 $\leq 4$
2 $\leq 19$
3 $\leq 46$
4 $\leq 85$
5 $\leq 136$
6 $\leq 199$
7 $\leq 274$
8 $\leq 361$
9 $\leq 460$
10 $\leq 571$
11 $\leq 694$
12 $\leq 829$
13 $\leq 976$
14 $\leq 1135$
15 $\leq 1306$
If $c(K)$, then $sm(K)$
$\geq 5$ $\geq 2$
$\geq 20$ $\geq 3$
$\geq 47$ $\geq 4$
$\geq 86$ $\geq 5$
$\geq 137$ $\geq 6$
$\geq 200$ $\geq 7$
$\geq 275$ $\geq 8$
$\geq 362$ $\geq 9$
$\geq 461$ $\geq 10$
$\geq 572$ $\geq 11$
$\geq 695$ $\geq 12$
$\geq 830$ $\geq 13$
$\geq 977$ $\geq 14$
$\geq 1136$ $\geq 15$
$\geq 1307$ $\geq 16$
Table 3: Several consequences of Corollary 5.8.1.
There are several more questions that spherical knot mosaics give rise to which the authors plan to explore in subsequent papers.
@@ Acknowledgments
The authors would like to thank Jennifer Schultens for her encouragement and advice throughout this project. We also thank Aaron Heap for his feedback and keen eye in catching some errors in the appendix of a previous version of this paper. The second author would like to thank Samantha Pezzimenti for introducing him to knot mosaics during the Unknot V conference at Seattle University and inspiring their work in this direction. Both authors thank Samantha Pezzimenti for her comments on and benignity of a previous version of this paper.
@@ Appendix A Planar Representation Examples
Planar representations of the first 35 unoriented prime knots as spherical knot mosaics. Due to Corollary 5.8.1, each of these planar representations realize their knot type’s spherical mosaic number, but in general these may not be minimally space efficient.
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]